# Derivative-Free Guidance in Continuous and Discrete Diffusion Models with Soft Value-Based Decoding

### v2

@@ 1 Introduction
The optimization of downstream reward functions using pre-trained diffusion models has been approached in various ways. In our work, we focus on non-fine-tuning-based methods because fine-tuning generative models (e.g., when using classifier-free guidance (Ho et al., 2020) or RL-based fine-tuning (Black et al., 2023; Fan et al., 2023; Uehara et al., 2024; Clark et al., 2023; Prabhudesai et al., 2023)) often becomes computationally intensive, especially as pre-trained generative models grow larger in the era of “foundation models”. Although classifier guidance and its variants (e.g., Dhariwal and Nichol (2021); Song et al. (2020); Chung et al. (2022); Bansal et al. (2023); Ho et al. (2022)) have shown success as non-fine-tuning methods in these settings, they face significant challenges. Firstly, as they would require constructing differentiable proxy models, they cannot directly incorporate useful domain-specific non-differentiable features (e.g., molecular/protein descriptors (van Westen et al., 2013; Ghiringhelli et al., 2015; Emonts and Buyel, 2023)) or non-differentiable reward feedback (e.g., physics-based simulations such as Vina and Rosetta (Trott and Olson, 2010; Alhossary et al., 2015; Alford et al., 2017)). This limitation also hinders the application of current classifier guidance methods to recently developed discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023) in a principled manner (i.e., without transforming the discrete space into the Euclidean space). Additionally, gradient computation can be memory-intensive with large proxy models.
@@
Table 1: A comparison of our method with a number of prior approaches. “Non-differentiable” refers to the method’s capability to operate without requiring differentiable proxy models. “No Training” means that no additional training of the diffusion model is required as long as we have access to the reward feedback.
@@ 2 Related Works
There are primarily two methods for optimizing downstream functions in diffusion models without fine-tuning. Regarding the comparison with the most related work (TDS), we defer the discussion to Section 5.4.
@@ 5.3 Extensions
5.4 Comparison between SVDD and Twisted Diffusion Sampler
The algorithm most closely related to SVDD, is the twisted diffusion sampler (TDS) (Wu et al., 2024; Trippe et al., 2022), which utilizes vanilla sequential Monte Carlo (Del Moral and Doucet, 2014; Kitagawa, 1993) in diffusion models. Although originally designed for conditioning, it can technically be applied to reward maximization. However, TDS requires resampling across the “entire” batch, complicating parallelization. Additionally, when batch sizes are small, as is often the case in recent large diffusion models, its performance may be suboptimal. In contrast, our method is easily parallelizable and can achieve strong performance even with a batch size of $1$, as resampling occurs for each individual sample. For further details, see Section A.
@@ Acknowledgement
Appendix A Difference between SVDD and Twisted Diffusion Sampler
Algorithm 4 TDS (for reward maximization)
@@ Appendix A Difference between SVDD and Twisted Diffusion Sampler
In this section, we compare our algorithm with TDS (Wu et al., 2024; Trippe et al., 2022), which uses sequential Monte Carlo for guidance. We then elaborate on the differences between SVDD and TDS. Notably, our SVDD is an instantiation of nested-IS SMC (Naesseth et al., 2019, Algorithm 5) in the literature on computational statistics, whereas TDS relies on vanilla sequential Monte Carlo.
A.1 Twisted Diffusion Sampler
@@
  • Wu et al. (2024) actually considers a scenario where the reward $r$ is a classifier. In Algorithm 4, we tailor the algorithm for reward maximization. Vice verisa, as we mentioned in Section 5.3, our SVDD can also operate effectively when $r$ is a classifier.
  • In Wu et al. (2024), the proposal distribution is not limited to the pre-trained model. Likewise, in our SVDD, we can select an arbitrary proposal distribution, as discussed in Section C.
@@
A.2 Comparison of TDS with SVDD
We now compare our SVDD with TDS. Here, we write a batch size of SVDD in $G$. Importantly, we note that our implementation is analogous to nested-IS SMC in the literature in computational statistics; hence, many differences between nested-IS SMC (Naesseth et al., 2019, 2015) and vanilla SMC are translated here.
We first reconsider the fundamental assumptions of each algorithm. SVDD’s performance, in terms of rewards, depends on the size of $M$ but is independent of the batch size $G$. In contrast, the performance of TDS depends on the batch size $N$. With this in mind, we compare the advantages of SVDD over TDS from various perspectives.
@@ A.2 Comparison of TDS with SVDD
SVDD is significantly easier to parallelize across multiple nodes. In contrast, TDS requires interaction between nodes for parallelization. This advantage is also well-documented in the context of nested-IS SMC versus vanilla SMC (Naesseth et al., 2019).
@@
Now, consider a scenario where the batch size is small, which often occurs due to memory constraints in large pre-trained diffusion models. In this case, while TDS may exhibit suboptimal performance, SVDD-PM can still achieve high performance by choosing a sufficiently large $M$ (assuming that models of the value functions are not significantly larger than those in pre-trained diffusion models).
In TDS, the “ratio” is approximated.
In SVDD, we approximate each $\exp(v_{t-1}(x_{t-1})/\alpha)$ as a weight. However, in TDS, the ratio is approximated as a weight:
@@
The key difference is that in TDS, both the numerator and the denominator are approximated, which could lead to greater error propagation.

### v3

@@ 1 Introduction
The optimization of downstream reward functions using pre-trained diffusion models has been approached in various ways. In our work, we focus on non-fine-tuning-based methods because fine-tuning generative models (e.g., when using classifier-free guidance (Ho et al., 2020) or RL-based fine-tuning (Black et al., 2023; Fan et al., 2023; Uehara et al., 2024; Clark et al., 2023; Prabhudesai et al., 2023)) often becomes computationally intensive, especially as pre-trained generative models grow larger in the era of “foundation models”. Although classifier guidance and its variants (e.g., Dhariwal and Nichol (2021); Song et al. (2020); Chung et al. (2022); Bansal et al. (2023); Ho et al. (2022)) have shown success as non-fine-tuning methods in these settings, they face significant challenges. Firstly, as they would require constructing differentiable proxy models, they cannot directly incorporate useful domain-specific non-differentiable features (e.g., molecular/protein descriptors (van Westen et al., 2013; Ghiringhelli et al., 2015; Gainza et al., 2020)) or non-differentiable reward feedback (e.g., physics-based simulations such as Vina and Rosetta (Trott and Olson, 2010; Alhossary et al., 2015; Alford et al., 2017)). This limitation also hinders the application of current classifier guidance methods to recently developed discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023) in a principled manner (i.e., without transforming the discrete space into the Euclidean space). Additionally, gradient computation can be memory-intensive with large proxy models.
@@
Table 1: A comparison of our method with a number of prior approaches. “Non-differentiable” refers to the method’s capability to operate without requiring differentiable proxy models. “No Training” means that no additional training of the diffusion model is required as long as we have access to the reward feedback. Regarding the comparison with standard SMC-based methods, refer to Section A.
@@ 2 Related Works
There are primarily two methods for optimizing downstream functions in diffusion models without fine-tuning. Regarding the comparison with the most related works based on sequential Monte Carlo (SMC), we defer the discussion to Section 5.4 and Section A.
@@ 5.3 Extensions
5.4 Comparison between SVDD and Standard SMC-Methods
The algorithm most closely related to SVDD, are several SMC-Methods for guidance (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023), which utilizes “standard” sequential Monte Carlo (Del Moral and Doucet, 2014; Kitagawa, 1993) in diffusion models. Although originally designed for conditioning, it can technically be applied to reward maximization. However, they require resampling across the “entire” batch, complicating parallelization. Additionally, when batch sizes are small, as is often the case in recent large diffusion models, its performance may be suboptimal. In contrast, our method is easily parallelizable and can achieve strong performance even with a batch size of $1$, as resampling occurs for each individual sample. For further details, see Section A.
@@ Acknowledgement
Appendix A Difference Between SVDD and “Standard” SMC-Methods
Algorithm 4 Guidance with “Standard” SMC (for reward maximization)
@@ Appendix A Difference Between SVDD and “Standard” SMC-Methods
In this section, we compare our algorithm with the SMC-based methods (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023) for guidance. While they originally aim to solve a conditioning problem, which is different from reward maximization, they algorithms can be converted to reward maximization. We then elaborate on the differences between SVDD, and them in the context of reward maximization. Notably, our SVDD is an instantiation of nested-IS SMC (Naesseth et al., 2019, Algorithm 5) in the literature on computational statistics, whereas these SMC-based Methods rely on standard sequential Monte Carlo.
A.1 SMC-Based Methods
@@
  • All of existing works Wu et al. (2024); Cardoso et al. (2023); Phillips et al. (2024); Dou and Song (2024) actually consider a scenario where the reward $r$ is a classifier. In Algorithm 4, we tailor the algorithm for reward maximization. Vice verisa, as we mentioned in Section 5.3, our SVDD can also operate effectively when $r$ is a classifier.
  • In Wu et al. (2024); Cardoso et al. (2023); Phillips et al. (2024), the proposal distribution is not limited to the pre-trained model. Likewise, in our SVDD, we can select an arbitrary proposal distribution, as discussed in Section C.
@@
A.2 Comparison of SVDD with “Standard” SMC-Based Methods (SSM) for Reward Maximization
We now compare our SVDD with “standard” SMC-Based Methods (SSM). Here, we write a batch size of SVDD in $G$. Importantly, we note that our implementation is analogous to nested-IS SMC in the literature in computational statistics; hence, many differences between nested-IS SMC (Naesseth et al., 2019, 2015) and pure SMC in computational statistics are translated here.
We first reconsider the fundamental assumptions of each algorithm. SVDD’s performance, in terms of rewards, depends on the size of $M$ but is independent of the batch size $G$. In contrast, the performance of SSM depends on the batch size $N$. With this in mind, we compare the advantages of SVDD over SSM from various perspectives.
@@ A.2 Comparison of SVDD with “Standard” SMC-Based Methods (SSM) for Reward Maximization
SVDD is significantly easier to parallelize across multiple nodes. In contrast, SSM requires interaction between nodes for parallelization. This advantage is also well-documented in the context of nested-IS SMC versus standard SMC (Naesseth et al., 2019).
@@
Now, consider a scenario where the batch size is small, which often occurs due to memory constraints in large pre-trained diffusion models. In this case, while SSM may exhibit suboptimal performance, SVDD-PM can still achieve high performance by choosing a sufficiently large $M$ (assuming that models of the value functions are not significantly larger than those in pre-trained diffusion models).
In SSM, the “ratio” is approximated.
In SVDD, we approximate each $\exp(v_{t-1}(x_{t-1})/\alpha)$ as a weight. However, in standard SMC, the ratio is approximated as a weight:
@@
The key difference is that in SSM, both the numerator and the denominator are approximated, which could lead to greater error propagation.